OpenStudy (anonymous):

Solve the equation 4sqrt/x-2=sqrt/5-x

OpenStudy (anonymous):

\[4 \sqrt{x-2} = \sqrt{5-x}\]Squaring both sides: 16 (x-2) = (5-x) 16x - 32 = 5 - x 17x = 37 x = 37/17

OpenStudy (anonymous):

ok